OpenStudy (anonymous):

Solve for x exactly (no graphing) in (ln x)^4 = ln x^4. You will get two answers. Their sum is approximately A) 4.291 B) 4.691 C) 5.091 D) 5.491 E) 5.891

OpenStudy (anonymous):

So if you just want a second opinion on your answer, what answer did you get?

OpenStudy (anonymous):

If you just want to get the answer, use wolframalpha. If you want to learn you'll need to put in some effort.

OpenStudy (anonymous):

Actually I have no idea how to solve this one. Im not sure what they want. Any ideas?

OpenStudy (anonymous):

Personally I would start with a substitution to make it easier to see what's going on

OpenStudy (anonymous):

For example, lets let k = ln(x)

OpenStudy (anonymous):

What would we have for our new version of the equation?

OpenStudy (anonymous):

(k)^4= k^4

OpenStudy (anonymous):

not quite.

OpenStudy (anonymous):

Thats basically the same thing on both sides of the equal sign

OpenStudy (anonymous):

It should be \( k^4 = ln(x^4)\)

OpenStudy (anonymous):

because in general for some value a, \[ln(a^4) \ne (ln\ a)^4\]

OpenStudy (anonymous):

But there is something we can do with the x^4 because of the properties of logarithms

OpenStudy (anonymous):

k^4= 4ln(x) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And since we said k = ln(x) ?

OpenStudy (anonymous):

k^4 = 4k?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

So solve for k

OpenStudy (anonymous):

do I divide by 4 first or do I take the 4th root?

OpenStudy (anonymous):

I would move it over to the same side so it equals 0.

OpenStudy (anonymous):

then factor it

OpenStudy (anonymous):

Actually that might not work out well

OpenStudy (anonymous):

I mean it's correct, but there might be something easier

OpenStudy (anonymous):

hmmmm im trying to factor it and im at k(k^3-4)

OpenStudy (anonymous):

Right, so either k = 0 or k^3 = 4

OpenStudy (anonymous):

That'll work ok

OpenStudy (anonymous):

So now lets go back to our definition of k

OpenStudy (anonymous):

ln(x) = 0 and ln(x)^3=4

OpenStudy (anonymous):

ln(x) = 0 & ln(x) = (cube root)4 ?

OpenStudy (anonymous):

Yep, now raise e to the power of both sides to get rid of the ln

OpenStudy (anonymous):

x= (cuberoot 4) e = 4.31?

OpenStudy (anonymous):

Should be: \[x = e^0 \text{ or } x = e^\sqrt[3]{4}\]

OpenStudy (anonymous):

Because we had \[ln(x) = 0 \text{ or } ln(x) = \sqrt[3]{4}\] and \[ln(a) = b \iff e^b = a\]

OpenStudy (anonymous):

5.891 is the answer. That was a tuffy for me.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Indeed. Takes a bit of thinking